# DNOI-4DRO: Deep 4D Radar Odometry with Differentiable Neural-Optimization Iterations

### v1

@@ 4.2 Supervision
4.3 Training Details
We train for 30 epochs on an A100 GPU with a batch size of 4. We use the Adam optimizer and start with an initial learning rate of $2e^{-4}$, which decays by a factor of 0.1 every 10 epochs. $N$ is set to be 512 in the proposed network. The radar point heights are constrained within the range of $[-2m,10m]$ to retain reliable points. During training, random transformation matrices are applied to the points and the ground truth pose to increase data diversity. We select 8 cluster centers. The iteration operator is unrolled for 15 iterations during training, with two bundle adjustment steps performed within each iteration. The ground truth pose of the first frame in each training sequence serves as the initial pose for all radar frames in that sequence, with the initial point motion flow set to zero.
4.4 4DRO System
@@ 4.3 Training Details
5 Experiments
We evaluate the method on two 4D radar odometry benchmarks, with dataset details in the appendix. We follow protocols of DVLO dvlo to evaluate our method with two metrics: (1) Average sequence translational RMSE (%). (2) Average sequence rotational RMSE (/100m).
5.1 Performance Evaluation
@@
CAO-RONet 0.049 0.034 0.042 0.040 0.064 0.089 0.082 0.063 0.135 0.076 0.102 0.013 0.024 0.045 0.071 0.051
Ours 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023$\downarrow$67.60.027$\downarrow$47.1%
Table 1: The 4D radar odometry experiment results on the VoD dataset vod . $t_{rel}$ and $r_{rel}$ denote the average translational RMSE (m/m) and rotational RMSE (/m), respectively, on all possible subsequences in the length of $20,40,...,160m$. All methods listed use 4D radar point clouds as input. Full A-LOAM is a complete SLAM system and others only include odometry.
@@
Ours R 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023$\downarrow$32.40.027$\downarrow$59.7%
Table 2: The experiment results on the VoD dataset vod using methods with LiDAR or 4D radar-camera inputs. ‘L’, ‘C’, and ‘R’ represent LiDAR, camera, and 4D radar, respectively.
Quantitative results are listed in Table 1. ICP-point2point (ICP-po2po) p2picp , ICP-point2plane (ICP-po2pl) p2plicp , GICP gicp , NDT ndt are several classical point cloud odometry methods. PWCLO-Net pwclo is currently the leading open-source deep learning-based LiDAR odometry method that directly processes raw 3D point clouds. A-LOAM loam is a conventional LiDAR odometry method that achieves state-of-the-art performance on the KITTI Odometry benchmark kitti . We use 4D radar point clouds as input for these methods. The experimental results indicate that, while these methods perform effectively on LiDAR point clouds, their performance deteriorates significantly on 4D radar point clouds due to the sparse, noisy, and non-panoramic characteristics of 4D radar data. RaFlow raflow and CMFlow cmflow are 4D radar-based scene flow estimation networks incorporating odometry estimation as an intermediate task. 4DRONet 4dronet and CAO-RONet cao are odometry networks specifically tailored for processing 4D radar data. Our method outperforms all previous 4D radar odometry methods due to its innovative architecture and robust point cloud feature extraction. Notably, our method reduces mean translation and rotation errors by 67.6% and 47.1%, respectively, compared to the previous best learning-based method.
Table 2 shows the evaluation results of Full A-LOAM and A-LOAM without mapping using 64-beam LiDAR point clouds as input, along with the results of the 4D radar-camera fusion odometry method (4DRVONet) 4drvonet . It can be seen that our method achieves competitive performance with Full A-LOAM in short-range localization. Compared with 4DRVONet, our method achieves a reduction of 72.3% in mean translation error and 59.7% in mean rotation error on the test sequences. The experiment results prove the effectiveness of our proposed neural-optimization iteration module and demonstrate its great potential in 4D radar odometry tasks.
@@
Table 3: The experiment results on the Snail-Radar dataset snailradar . $t_{rel}$ and $r_{rel}$ denote the average translational RMSE (%) and rotational RMSE (/100m), respectively, on all possible subsequences in the length of $100,200,...,800m$. ‘L’ and ‘R’ represent LiDAR and 4D radar, respectively.
Compared with the VoD dataset vod , the Snail-Radar Dataset snailradar contains longer trajectories and more complex environments, posing greater challenges to odometry stability. We compare our method with ICP-based point cloud odometry methods and 4DRadarSLAM 4dradarslam , a classical 4D radar SLAM method. We turn off the loop closure detection in 4DRadarSLAM to ensure a fair comparison. Additionally, we present the results of Full A-LOAM and A-LOAM without mapping, using 32-beam LiDAR point clouds as input. Table 3 indicates that, while our method is slightly less effective than Full A-LOAM in long-range odometry localization, it narrows the performance gap with leading LiDAR odometry methods and significantly outperforms A-LOAM without mapping. Furthermore, compared with 4DRadarSLAM, our method reduces mean translation and rotation errors by 65.7% and 71.6%, respectively.
5.2 Ablation Study
@@
Table 4: The ablation study results of 4D Radar odometry on the VoD dataset vod .
We perform ablation experiments on the Snail-Radar Dataset to evaluate the effectiveness of each module in DNOI-4DRO. The training/testing details are the same as in Sec. 4.
@@
5.3 Visualization
Figure 4: 3D and 2D trajectory results for VoD test sequences 19 (a) and 22 (b), and Snail-Radar test sequences st (c) and (d). Our method obtains the most accurate trajectory.
Figure 5: The Visualization of clustering results.
@@ 5.3 Visualization
We visualize 3D and 2D trajectories based on our estimated pose on the VoD and Snail-Radar datasets, as shown in Fig. 4. These figures show that our odometry can track the trajectory of the ground truth fairly well. Although we do not have the mapping procedure, our odometry achieves trajectory accuracy comparable to that of Full A-LOAM in short-range localization. Furthermore, our approach significantly reduces odometry drift along the Z-axis compared to other methods. Fig. 5 illustrates the clustering results. The clustering-based method effectively mitigates the constraints of the receptive field, facilitating extracting features that belong to the same category (e.g., vehicles, curbs, walls). This approach enables point features to exhibit class awareness, improving their representational capacity.
6 Conclusion and Limitation
@@
We introduce DNOI-4DRO, an end-to-end neural architecture for 4D radar odometry, which combines the strengths of both classical approaches and deep networks through a differentiable neural-optimization iteration module. We compare against existing approaches and show strong performance across several datasets. The current research primarily focuses on optimization techniques based on sliding windows. Effective global optimization can significantly enhance pose estimation accuracy. Future efforts will be directed toward exploring more efficient strategies for keyframe selection and the development of global optimization frameworks.
Appendix A Technical Appendices and Supplementary Material
A.1 Overview
This supplementary material provides more detailed implementation specifics in Sec. A.2. Sec. A.3 presents experimental results regarding runtime. Sec. A.4 shows additional qualitative evaluation results. Sec. A.5 discusses the broader impacts of the algorithm introduced in this paper.
A.2 Detailed Implementation Specifics
A.2.1 Data Augmentation Parameters
We adopt the data augmentation method from [6]. Specifically, we expand the training dataset by generating augmentation matrix $\mathbf{T}^{\text{aug}}$ through rotation matrix $\mathbf{R}^{\text{aug}}$ and translation vector $\mathbf{t}^{\text{aug}}$. Varied values of the yaw-pitch-roll Euler angles are generated by Gaussian distribution around $0^{\circ}$. We set the standard deviations of the yaw, pitch, and roll Euler angles to $0.05^{\circ}$, $0.01^{\circ}$, and $0.01^{\circ}$, respectively. We set the standard deviations for the translation in the XYZ directions to $0.2m$, $1.0m$, and $0.05m$, respectively. In these Gaussian distributions, we select the data in the range of 2 times standard deviation around the mean value for data augmentation.
The composed $\mathbf{T}^{\text{aug}}$ from $\mathbf{R}^{\text{aug}}$ and $\mathbf{t}^{\text{aug}}$ is then used to augment the $\mathbf{PC}^{2}$ to obtain new point clouds $\mathbf{PC}^{2,\text{aug}}$.
A.2.2 The Detailed Introduction of the Dataset
View-of-Delft Automotive Dataset (VoD) [25] is collected from the Delft campus, suburbs, and old town, integrating data from various sensors, including 4D radar. The dataset comprises 8682 frames, organized into 24 sequences based on frame continuity. We adopt the predefined sets from CMFlow [34] to evaluate the network’s performance for training, validation, and testing purposes.
Snail-Radar Dataset [26] is designed explicitly for 4D radar-based localization and mapping, covering campus and highway scenes under varied weather and lighting conditions. The dataset comprises 44 sequences collected across eight distinct scenes. One sequence is selected from each scene to avoid trajectory redundancy, resulting in eight sequences. Among these, five sequences (52,540 frames) are allocated for model training, while the remaining three sequences (23,037 frames) are designated for testing. The sequence lengths and weather and lighting conditions during data collection are presented in Table 5. As shown in Fig. 6, all test sequences were collected on rainy nights, where adverse weather conditions significantly impacted odometry stability. However, despite the challenging weather conditions, the method proposed in this paper still achieves high accuracy (Table 3), further confirming the approach’s effectiveness and the potential of 4D radar for application in complex environments.
Scene Sequence Split Dist.(m)/Dur.(sec) Weather/Lighting
bc 20231105/6 train 251/144 moderate, rain
sl 20231019/1 train 1919/460 night
ss 20231105/4 train 895/395 heavy rain
iaef 20240115/2 train 6641/1374
81r 20240116 eve/3 train 8521/1293 night
if 20231213/4 test 2227/494 light rain, night
iaf 20231213/2 test 4603/938 light rain, night
st 20231213/1 test 276/126 light rain, night
Table 5: Description of Snail-Radar Dataset [26].
Figure 6: Examples of three test sequences from the Snail-Radar dataset [26], all captured in nighttime rainy conditions.
A.3 Runtime Analysis
Module Feature Extraction Update Operator Total
Runtime(ms) 6.8 36.4 43.2
Table 6: Average inference time of different modules on the test sequence of the VoD dataset[25].
Efficiency is another highly significant factor in real-time SLAM systems. We evaluate the runtime of each module in DNOI-4DRO on the VoD dataset, as shown in Table 6. Since the 4D radar points in the VoD dataset are captured at a 10Hz frequency, our method has only 43.2ms inference time, which has the potential for real-time application.
A.4 Additional Qualitative Results
A.4.1 Confidence Weight Visualization
Figure 7: Confidence weight visualization. (a) Confidence in the X direction. (b) Confidence in the Y direction.
Fig. 7 shows the confidence weights estimated by the network on the VoD dataset [25]. The figure demonstrates that the network assigns varying confidence weights to the same point across different directions. In the X direction, high-confidence points are primarily selected from the vertical plane at intermediate distances, as these points provide stable longitudinal constraints. For the Y direction, high-confidence points are concentrated on those located along the road edge and the left-right planes, which offer reliable lateral constraints. Additionally, points on dynamic objects, such as cyclists, are assigned low confidence, thereby minimizing their adverse impact on the odometry estimation.
A.4.2 Average Translation and Rotation Errors
[…]
Fig. 8(a) shows the average translation and rotation error on VoD test sequences. Our method performs better in translation and rotation errors across various path lengths. Notably, our method exhibits a clear advantage compared to other odometry methods that use 4D radar point clouds as input. Additionally, our method yields superior results in most path lengths compared to A-LOAM [46] with mapping optimization, which uses LiDAR point clouds as input.
Fig. 8(b) presents the average translation and rotation error on the Snail-Radar test sequences. The longer trajectories on the Snail-Radar dataset allow the mapping optimization in Full A-LOAM to demonstrate its advantages fully. Although our method slightly underperforms compared to Full A-LOAM, which utilizes 32-beam LiDAR point clouds as input, it substantially narrows the performance gap with Full A-LOAM compared to other methods. Furthermore, considering that the number of 4D radar points is approximately 1% of the 32-beam LiDAR point clouds and that our method does not include a mapping optimization process, we believe the performance gap with Full A-LOAM is within an acceptable range, with potential for further enhancement.
A.4.3 More Trajectory Results on VoD Dataset
Figure 9: 2D trajectory results on VoD test sequences with ground truth. Our method achieves trajectory accuracy comparable to that of Full A-LOAM. (a) 2D Trajectory Plots of Seq. 03. (b) 2D Trajectory Plots of Seq. 04. (c) 2D Trajectory Plots of Seq. 09. (c) 2D Trajectory Plots of Seq. 24.
We list all visualized trajectory results on test sequences of VoD dataset [25] with the ground truth in Fig. 9, excluding the results in our submitted paper. The results demonstrate that our method outperforms existing 4D radar odometry methods and the A-LOAM without mapping using 64-beam LiDAR point clouds across all test sequences. Furthermore, it exceeds the performance of both the 4D radar-camera fusion odometry method and the Full A-LOAM using 64-beam LiDAR point clouds in most sequences.
A.4.4 Reconstruction
Figure 10: Example reconstruction on VoD dataset [25].
A.5 Broder Impacts
4D millimeter-wave radar is a low-cost, all-weather sensor. The 4D radar-based odometry method proposed in this paper enables high-precision localization for autonomous driving and robotics across diverse environments, particularly in challenging weather conditions. Moreover, the method demonstrates strong potential for real-time applications, facilitating its deployment in practical scenarios. By offering an efficient and robust solution, it strongly supports technological advancements in autonomous driving and robotics.

### v2

@@ Introduction
Training Details
We train for 30 epochs on an A100 GPU with a batch size of 4. We use the Adam optimizer and start with an initial learning rate of $2e^{-4}$, which decays by a factor of 0.1 every 10 epochs. $N$ is set to be 512 in the proposed network. The radar point heights are constrained within the range of [2m,10m][-2m,10m] to retain reliable points. During training, random transformation matrices are applied to the points and the ground truth pose to increase data diversity. We select 8 cluster centers. The iteration operator is unrolled for 15 iterations during training, with two bundle adjustment steps performed within each iteration. The ground truth pose of the first frame in each training sequence serves as the initial pose for all radar frames in that sequence, with the initial point motion flow set to zero.
4DRO System
@@
Experiments
We evaluate the method on two 4D radar odometry benchmarks, with dataset details in the appendix. We follow protocols of DVLO (dvlo) to evaluate our method with two metrics: (1) Average sequence translational RMSE (%). (2) Average sequence rotational RMSE (/100m).
Performance Evaluation
@@
CAO-RONet 0.049 0.034 0.042 0.040 0.064 0.089 0.102 0.013 0.024 0.045 0.082 0.063 0.135 0.076 0.071 0.051
Ours 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023($\downarrow$67.6%) 0.027($\downarrow$47.1%)
Table 1: The 4D radar odometry experiment results on the VoD dataset (vod). $t_{rel}$ and $r_{rel}$ denote the average translational RMSE (m/m) and rotational RMSE (/m), respectively, on all possible subsequences in the length of 20,40,,160m20,40,...,160m. All methods listed use 4D radar point clouds as input. Full A-LOAM is a complete SLAM system and others only include odometry. The best results are bold.
@@
Ours R 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023($\downarrow$32.4%) 0.027($\downarrow$59.7%)
Table 2: The experiment results on the VoD dataset (vod) using methods with LiDAR or 4D radar-camera inputs. ‘L’, ‘C’, and ‘R’ represent LiDAR, camera, and 4D radar, respectively. The best results are bold.
Quantitative results are listed in Table 1. ICP-point2point (ICP-po2po) (p2picp), ICP-point2plane (ICP-po2pl) (p2plicp), GICP (gicp), NDT (ndt) are several classical point cloud odometry methods. TransLO (translo) and EfficientLO (efficientlo) rely on the imaging principle of LiDAR to project point clouds onto a 2D cylindrical surface as network input. However, 4D radar operates with a different imaging mechanism and produces only about 512 sparse points per frame, making projection-based approaches unsuitable for such data. Therefore, we primarily compare our method with leading open-source LiDAR odometry techniques that directly process raw 3D point clouds, such as PWCLO-Net (pwclo). A-LOAM (loam) is a conventional LiDAR odometry method that achieves state-of-the-art performance on the KITTI Odometry benchmark (kitti). We use 4D radar point clouds as input for these methods. The experimental results indicate that, while these methods perform effectively on LiDAR point clouds, their performance deteriorates significantly on 4D radar point clouds due to the sparse, noisy, and non-panoramic characteristics of 4D radar data. RaFlow (raflow) and CMFlow (cmflow) are 4D radar-based scene flow estimation networks incorporating odometry estimation as an intermediate task. 4DRONet (4dronet) and CAO-RONet (cao) are odometry networks specifically tailored for processing 4D radar data. Our method outperforms all previous 4D radar odometry methods due to its innovative architecture and robust point cloud feature extraction. Notably, our method reduces mean translation and rotation errors by 67.6% and 47.1%, respectively, compared to the previous best learning-based method.
Table 2 shows the evaluation results of Full A-LOAM and A-LOAM without mapping using 64-beam LiDAR point clouds as input, along with the results of the 4D radar-camera fusion odometry method (4DRVONet) (4drvonet). It can be seen that our method achieves competitive performance with Full A-LOAM in short-range localization. Compared with 4DRVONet, our method achieves a reduction of 72.3% in mean translation error and 59.7% in mean rotation error on the test sequences. The experiment results prove the effectiveness of our proposed neural-optimization iteration module and demonstrate its great potential in 4D radar odometry tasks.
@@
Table 3: The experiment results on the Snail-Radar dataset (snailradar). $t_{rel}$ and $r_{rel}$ denote the average translational RMSE (%) and rotational RMSE (/100m), respectively, on all possible subsequences in the length of 100,200,,800m100,200,...,800m. ‘L’ and ‘R’ represent LiDAR and 4D radar, respectively. The best results are shown in bold, and the second-best results are underlined.
Compared with the VoD dataset (vod), the Snail-Radar Dataset (snailradar) contains longer trajectories and more complex environments, posing greater challenges to odometry stability. We compare our method with ICP-based point cloud odometry methods and 4DRadarSLAM (4dradarslam), a classical 4D radar SLAM method. We turn off the loop closure detection in 4DRadarSLAM to ensure a fair comparison. Additionally, we present the results of Full A-LOAM and A-LOAM without mapping, using 32-beam LiDAR point clouds as input. Table 3 indicates that, while our method is slightly less effective than Full A-LOAM in long-range odometry localization, it narrows the performance gap with leading LiDAR odometry methods and significantly outperforms A-LOAM without mapping. Furthermore, compared with 4DRadarSLAM, our method reduces mean translation and rotation errors by 65.7% and 71.6%, respectively.
Ablation Study
@@
Table 4: The ablation study results of 4D Radar odometry on the VoD dataset (vod). The best results are bold.
Figure 4: 3D and 2D trajectory results for VoD test sequences 19 and $2$, and Snail-Radar test sequences st. Our method obtains the most accurate trajectory.
Figure 5: Visualization of clustering results. 4D radar points are color-coded based on their assigned categories.
We perform ablation experiments on the VoD Dataset to evaluate the effectiveness of each module in DNOI-4DRO.
@@
Visualization
@@
We visualize 3D and 2D trajectories based on our estimated pose on the VoD and Snail-Radar datasets, as shown in Figure 4. These figures show that our odometry can track the trajectory of the ground truth fairly well. Although we do not have the mapping procedure, our odometry achieves trajectory accuracy comparable to that of Full A-LOAM in short-range localization. Furthermore, our approach significantly reduces odometry drift along the Z-axis compared to other methods. Figure 5 illustrates the clustering results. The clustering-based method effectively mitigates the constraints of the receptive field, facilitating extracting features that belong to the same category (e.g., vehicles, curbs, walls). This approach enables point features to exhibit class awareness, improving their representational capacity.
Conclusion
@@
We introduce DNOI-4DRO, an end-to-end neural architecture for 4D radar odometry, which combines the strengths of both classical approaches and deep networks through a differentiable neural-optimization iteration module. We compare against existing approaches and show strong performance across several datasets.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant 52325212.